# VLM-Pruner: Buffering for Spatial Sparsity in an Efficient VLM Centrifugal Token Pruning Paradigm

### v2

@@ 1 Introduction
Building on the powerful reasoning capabilities of large language models (LLMs) [gpt4, vicuna, llama], vision–language models (VLMs) [llava, llava15, llavanext, llavavideo, qwen2vl] integrate a visual encoder (e.g., ViT [vit]) to introduce visual modalities. This design enables VLMs to excel in tasks such as image captioning, visual question answering, video understanding, and multimodal reasoning. This integration significantly broadens their range of practical applications. However, the visual modality inherently carries high information density, particularly in high-resolution images [llavanext] and multi-frame videos [llavavideo], which results in a number of visual tokens that can exceed textual tokens by hundreds or even thousands of times. When these abundant visual tokens are merely projected and concatenated with textual tokens before being fed into the decoder layers of LLMs [llava], where attention computation exhibits quadratic complexity, the overall efficiency is substantially degraded.
@@
As discussed above, the excessive number of visual tokens not only reduces computational efficiency but also introduces significant redundancy: 1) Cross-modal redundancy, where only visual regions relevant to the textual query are necessary; and 2) Intra-modal redundancy, where tokens representing similar regions of the same object carry overlapping information. Recent studies have addressed these issues through training-free token pruning methods that select a representative subset of visual tokens, typically focusing on two strategies: 1) importance-based pruning [fastv, hired, pyramiddrop, sparsevlm], which leverages intra- or cross-modal attention scores to retain salient tokens; and 2) redundancy-based pruning [dart, divprune], which employs greedy algorithms to iteratively choose tokens with low similarity to those selected.
@@
However, the above strategies still perform suboptimally and exhibit notable drawbacks: 1) The importance-driven strategies focus on similar local regions often leads to redundancy. Methods based on visual importance like FastV [fastv] and on visual-language importance like SparseVLM [sparsevlm] often perform no better than random token pruning. In some cases they even perform worse [dart], because importance driven criteria tend to preserve multiple similar local regions around the same focused object, causing substantial redundancy. Although methods such as FiCoCo [ficoco] and MustDrop [mustdrop] apply localized penalties to prevent redundancy from overly clustered tokens, as shown in the person of Case 2 in Fig. 2, moderate clustering can, in fact, better preserve local details.; 2) The redundancy-reduction strategies leads to dispersed token selections. Methods targeting visual redundancy like DivPrune [divprune] and those addressing cross-modal redundancy like DART [dart] generate token selections that are dispersed yet incomplete, failing to capture the fine-grained details of objects. As shown in Fig. 2, edge background regions have low similarity to the main object, so DivPrune and DART are prone to selecting edge background tokens and jumping erratically between background and foreground, resulting in a scattered selection.
@@
Accordingly, we propose a training-free centrifugal token pruning paradigm, termed VLM-Pruner, that effectively balances token redundancy and local-detail completeness. VLM-Pruner follows a near-to-far token selection process, starting with pivot tokens, then prioritizing spatially adjacent tokens, and finally recovering from discarded tokens. This ensures an orderly sequence and avoids the chaotic, incomplete coverage from arbitrary selection. For example, in the first case of Fig. 2, four tokens closely aligned in order on the fork at the left of the image are selected when compared to other baselines; in the third case, VLM-Pruner captures comprehensive truck details, including the body, tires, and windshield. Specifically, a minimal pivot token set with maximum minimum distance [maxmin] is firstly identified to coarsely represent distinct subjects. Secondly, the Buffering for Spatial Sparsity (BSS) criterion leverages the minimum spatial distance between candidate tokens and the selected set to preferentially select spatially adjacent tokens with low redundancy, thereby ensuring an orderly selection process. Tokens are then greedily chosen in ascending similarity order, with parallel processing for acceleration. Finally, the remaining tokens are matched to the selected set based on maximum similarity and merged through Similarity-Weighted Aggregation (SWA) to recover the outermost information. In summary, our major contributions are as follows:
@@ 2.1 Efficient Vision-Language Models
Vision–Language Models (VLMs) [vit] align visual encoders with large language models to support captioning, VQA, grounding, and reasoning, with image-based [llava, llava15, llavanext, qwen2vl] and video-based [llavavideo] variants. A common recipe couples LLMs, which perform strong at text processing, with high-capacity visual encoders such as ViT [vit], CLIP-ViT [clipvit], or SigLIP [siglip] via cross-modal alignment.
Efficiency efforts aim to preserve multimodal capability while cutting compute, memory, and latency. (i) Connector-tuning freezes the backbones and trains lightweight adapters, e.g., BLIP-2 [blip], avoiding full end-to-end retraining. (ii) Content-adaptive budgeting allocates resolution and token budgets based on input saliency, as in Qwen2-VL [qwen2vl]. (iii) Distill-and-prune compression transfers knowledge [kd] and removes modality-aware redundancy. EfficientVLM [efficientvlm] retains 44.3% parameters and yields the inference speedup with minimal quality loss. Expert Merging [zhang2025expert] integrates capabilities of multiple VLMs.
@@ 2.2 Token Pruning
Token pruning reduces the quadratic cost of attention by compressing tokens at their source. ToMe [tome] pioneered progressive token merging for ViTs [vit], inspiring VLM pruning methods that are either training-required or training-free. Among training-required approaches, ATP-LLaVA [atpllava] learns layer-wise, instance-specific pruning ratios via MLPs, but adds training overhead. We focus on training-free pruning, which falls into importance-based and redundancy-based families.
Importance-driven methods keep tokens with high (self- or cross-modal) attention: FastV [fastv] uses self-attention on visual tokens; SparseVLM [sparsevlm] jointly prunes visual/text tokens via cross-modal attention; HiRED [hired] and MustDrop [mustdrop] favor tokens strongly attending to [CLS]; PDrop [pyramiddrop] retains those attentive to the last-instruction token. Redundancy-reduction methods prefer diversity: DART [dart] incrementally selects tokens with low similarity to text–visual anchors, while DivPrune [divprune] greedily preserves tokens least similar to those already kept.
@@ 3.2 Buffering for Spatial Sparsity (BSS) Criterion
When $\lambda=0$, the objective $\mathcal{F}$ reduces to a standard monotone submodular function [cost], for which the greedy algorithm achieves the classical approximation guarantee [greedy]. When $\lambda>0$, as the iterations proceed, the number of tokens in the selected set $\mathcal{S}$ gradually increases, causing the remaining candidate tokens to lie closer to the already selected ones in spatial space. Consequently, the normalized spatial distance term $\bar{\delta}_{i}(S)$ in Eq. (5) continuously decreases with each iteration. This monotonic decay preserves the overall submodular structure of $\mathcal{F}$ across iterations and therefore still permits an efficient greedy optimization scheme. Meanwhile, for $\lambda>0$, tokens that are spatially farther from the already selected set are assigned a larger scaling coefficient in their similarity computation, making them less likely to be selected in subsequent iterations. As shown in Fig. 3, this adaptive weighting mechanism effectively realizes the BSS effect, ensuring that nearby tokens are prioritized first, thereby promoting dense local coverage before the selection expands outward.
@@ 3.3.1 Stage 1: Pivot Initialization
We denote $\kappa$ as the number of initial pivots. The token keyskeys inherently summarize the semantic characteristics of each token [attention] and contain less redundant information due to their lower dimensionality. Therefore, we select $\kappa$ diverse pivots $\mathcal{S}_{\kappa}$ in the token keyskeys space using a max-min strategy:
@@ 4.1 Benchmarks and Base VLMs
To comprehensively evaluate the capability of our proposed VLM-Pruner, which is aimed at multimodal reasoning and understanding, we conduct extensive experiments on a diverse set of benchmark datasets and VLMs with varying architectures. Specifically, we select a broad range of commonly used tasks and benchmarks covering multiple-choice question answering (QA), optical character recognition (OCR), and open-ended QA. The evaluation includes 9 image-language benchmarks, namely GQA [gqa], MMB [mmbench], MME [mme], POPE [pope], SQA [sqa], TextVQA [textvqa], OCRBench [ocrbench], SEEDBench [seed], and OKVQA [okvqa], as well as 4 video-language benchmarks, namely EgoSchema [egoschema], NExTQA [nextqa], VideoMME [videomme], and EgoPlan [egoplan]. For base VLMs, we choose image-based models including LLaVA-1.5-7B [llava15], LLaVA-1.5-13B [llava15], LLaVA-Next-7B [llavanext], and Qwen2-VL-7B [qwen2vl], together with the video-based model LLaVA-Video-7B [llavavideo]. Please refer to the supplementary materials for more details.
@@ 4.2 Implementation Details
For LLaVA-1.5 [llava] with a fixed number of visual tokens, we set $H=W=24$ ($N=576$), so that $D_{\max}=24\sqrt{2}$, which exactly matches the normalization in Eq. (5). For LLaVA-Next [llavanext] and Qwen2-VL [qwen2vl], which adopt a dynamic number of tokens depending on the image resolution, $D_{\max}$ is computed according to the actual height and width of the token feature map. For LLaVA-Video [llavavideo], which involves three spatial-temporal dimensions (height, width, and frames), the computation is correspondingly performed in the 3D coordinate space. Meanwhile, we set $\lambda=0.5$, $\tau^{(0)}=0.8$, $\Delta\tau=0.1$, $q=256$, $\beta=0.3$, and $B=16$.
All experiments are performed on a setup consisting of 8 GPUs, each with 32 GB of VRAM. We adhere to the official evaluation protocols established for both LLaVA-1.5 and LLaVA-Next. Additionally, for evaluating Qwen2-VL and LLaVA-Video, we utilize the lmms-evals package [lmmseval]. All results are obtained with a batch size of 1.
@@ 4.3 Comparative Experiment
We select methods based on importance-based baselines, such as FastV [fastv], SparseVLM [sparsevlm], PDrop [pyramiddrop], and MustDrop [mustdrop], as well as methods based on redundancy-based baselines, such as DART [dart] and DivPrune [divprune], as the comprehensive baselines for comparison.
@@ 4.4 Efficiency Analysis
We analyze the efficiency of VLM-Pruner on LLaVA-1.5-7B and Qwen2-VL-7B. Specifically, we report FLOPs following [fastv] and total inference time on the POPE and OCRBench benchmarks under a $\downarrow$88.9% pruning rate (i.e., retaining 11.1% of visual tokens), as shown in Table 5. VLM-Pruner outperforms the importance-based baselines SparseVLM and MustDrop in both speedup and FLOPs. In comparisons with redundancy-based methods, DivPrune exhibits slightly fewer FLOPs. This is because it prunes at layer 0, while VLM-Pruner and DART start at layer 2. However, this DivPrune setting comes at the cost of significantly poorer performance compared to our method (Sec. 4.3). VLM-Pruner’s inference speed matches DART. On Qwen2-VL-7B, which has more raw visual tokens, VLM-Pruner excels in detail-intensive tasks like OCRBench. It achieves the fastest speed (1.60×) and the best performance (Table 3). These results demonstrate VLM-Pruner’s superior capability for handling high-resolution inputs.
@@ 7 Details for Benchmarks
We select a broad range of commonly used tasks and benchmarks. Specifically, the evaluation includes 9 image-language benchmarks, namely GQA [gqa], MMB [mmbench], MME [mme], POPE [pope], SQA [sqa], TextVQA [textvqa], OCRBench [ocrbench], SEEDBench [seed], and OKVQA [okvqa], as well as 4 video-language benchmarks, namely EgoSchema [egoschema], NExTQA [nextqa], VideoMME [videomme], and EgoPlan [egoplan].

### v3

@@ 1 Introduction
Building on the powerful reasoning capabilities of large language models (LLMs) [1, 8, 31], vision–language models (VLMs) [20, 18, 19, 43, 33] integrate a visual encoder (e.g., ViT [9]) to introduce visual modalities. This design enables VLMs to excel in tasks such as image captioning, visual question answering, video understanding, and multimodal reasoning. This integration significantly broadens their range of practical applications. However, the visual modality inherently carries high information density, particularly in high-resolution images [19] and multi-frame videos [43], which results in a number of visual tokens that can exceed textual tokens by hundreds or even thousands of times. When these abundant visual tokens are merely projected and concatenated with textual tokens before being fed into the decoder layers of LLMs [20], where attention computation exhibits quadratic complexity, the overall efficiency is substantially degraded.
@@
As discussed above, the excessive number of visual tokens not only reduces computational efficiency but also introduces significant redundancy: 1) Cross-modal redundancy, where only visual regions relevant to the textual query are necessary; and 2) Intra-modal redundancy, where tokens representing similar regions of the same object carry overlapping information. Recent studies have addressed these issues through training-free token pruning methods that select a representative subset of visual tokens, typically focusing on two strategies: 1) importance-based pruning [6, 3, 37, 42], which leverages intra- or cross-modal attention scores to retain salient tokens; and 2) redundancy-based pruning [35, 2], which employs greedy algorithms to iteratively choose tokens with low similarity to those selected.
@@
However, the above strategies still perform suboptimally and exhibit notable drawbacks: 1) The importance-driven strategies focus on similar local regions often leads to redundancy. Methods based on visual importance like FastV [6] and on visual-language importance like SparseVLM [42] often perform no better than random token pruning. In some cases they even perform worse [35], because importance driven criteria tend to preserve multiple similar local regions around the same focused object, causing substantial redundancy. Although methods such as FiCoCo [11] and MustDrop [21] apply localized penalties to prevent redundancy from overly clustered tokens, as shown in the person of Case 2 in Fig. 2, moderate clustering can, in fact, better preserve local details.; 2) The redundancy-reduction strategies leads to dispersed token selections. Methods targeting visual redundancy like DivPrune [2] and those addressing cross-modal redundancy like DART [35] generate token selections that are dispersed yet incomplete, failing to capture the fine-grained details of objects. As shown in Fig. 2, edge background regions have low similarity to the main object, so DivPrune and DART are prone to selecting edge background tokens and jumping erratically between background and foreground, resulting in a scattered selection.
@@
Accordingly, we propose a training-free centrifugal token pruning paradigm, termed VLM-Pruner, that effectively balances token redundancy and local-detail completeness. VLM-Pruner follows a near-to-far token selection process, starting with pivot tokens, then prioritizing spatially adjacent tokens, and finally recovering from discarded tokens. This ensures an orderly sequence and avoids the chaotic, incomplete coverage from arbitrary selection. For example, in the first case of Fig. 2, four tokens closely aligned in order on the fork at the left of the image are selected when compared to other baselines; in the third case, VLM-Pruner captures comprehensive truck details, including the body, tires, and windshield. Specifically, a minimal pivot token set with maximum minimum distance [29] is firstly identified to coarsely represent distinct subjects. Secondly, the Buffering for Spatial Sparsity (BSS) criterion leverages the minimum spatial distance between candidate tokens and the selected set to preferentially select spatially adjacent tokens with low redundancy, thereby ensuring an orderly selection process. Tokens are then greedily chosen in ascending similarity order, with parallel processing for acceleration. Finally, the remaining tokens are matched to the selected set based on maximum similarity and merged through Similarity-Weighted Aggregation (SWA) to recover the outermost information. In summary, our major contributions are as follows:
@@ 2.1 Efficient Vision-Language Models
Vision–Language Models (VLMs) [9] align visual encoders with large language models to support captioning, VQA, grounding, and reasoning, with image-based [20, 18, 19, 33] and video-based [43] variants. A common recipe couples LLMs, which perform strong at text processing, with high-capacity visual encoders such as ViT [9], CLIP-ViT [28], or SigLIP [39] via cross-modal alignment.
Efficiency efforts aim to preserve multimodal capability while cutting compute, memory, and latency. (i) Connector-tuning freezes the backbones and trains lightweight adapters, e.g., BLIP-2 [16], avoiding full end-to-end retraining. (ii) Content-adaptive budgeting allocates resolution and token budgets based on input saliency, as in Qwen2-VL [33]. (iii) Distill-and-prune compression transfers knowledge [12] and removes modality-aware redundancy. EfficientVLM [34] retains 44.3% parameters and yields the inference speedup with minimal quality loss. Expert Merging [40] integrates capabilities of multiple VLMs.
@@ 2.2 Token Pruning
Token pruning reduces the quadratic cost of attention by compressing tokens at their source. ToMe [4] pioneered progressive token merging for ViTs [9], inspiring VLM pruning methods that are either training-required or training-free. Among training-required approaches, ATP-LLaVA [38] learns layer-wise, instance-specific pruning ratios via MLPs, but adds training overhead. We focus on training-free pruning, which falls into importance-based and redundancy-based families.
Importance-driven methods keep tokens with high (self- or cross-modal) attention: FastV [6] uses self-attention on visual tokens; SparseVLM [42] jointly prunes visual/text tokens via cross-modal attention; HiRED [3] and MustDrop [21] favor tokens strongly attending to [CLS]; PDrop [37] retains those attentive to the last-instruction token. Redundancy-reduction methods prefer diversity: DART [35] incrementally selects tokens with low similarity to text–visual anchors, while DivPrune [2] greedily preserves tokens least similar to those already kept.
@@ 3.2 Buffering for Spatial Sparsity (BSS) Criterion
When $\lambda=0$, the objective $\mathcal{F}$ reduces to a standard monotone submodular function [14], for which the greedy algorithm achieves the classical approximation guarantee [27]. When $\lambda>0$, as the iterations proceed, the number of tokens in the selected set $\mathcal{S}$ gradually increases, causing the remaining candidate tokens to lie closer to the already selected ones in spatial space. Consequently, the normalized spatial distance term $\bar{\delta}_{i}(S)$ in Eq. (5) continuously decreases with each iteration. This monotonic decay preserves the overall submodular structure of $\mathcal{F}$ across iterations and therefore still permits an efficient greedy optimization scheme. Meanwhile, for $\lambda>0$, tokens that are spatially farther from the already selected set are assigned a larger scaling coefficient in their similarity computation, making them less likely to be selected in subsequent iterations. As shown in Fig. 3, this adaptive weighting mechanism effectively realizes the BSS effect, ensuring that nearby tokens are prioritized first, thereby promoting dense local coverage before the selection expands outward.
@@ 3.3.1 Stage 1: Pivot Initialization
We denote $\kappa$ as the number of initial pivots. The token keyskeys inherently summarize the semantic characteristics of each token [32] and contain less redundant information due to their lower dimensionality. Therefore, we select $\kappa$ diverse pivots $\mathcal{S}_{\kappa}$ in the token keyskeys space using a max-min strategy:
@@ 4.1 Benchmarks and Base VLMs
To comprehensively evaluate the capability of our proposed VLM-Pruner, which is aimed at multimodal reasoning and understanding, we conduct extensive experiments on a diverse set of benchmark datasets and VLMs with varying architectures. Specifically, we select a broad range of commonly used tasks and benchmarks covering multiple-choice question answering (QA), optical character recognition (OCR), and open-ended QA. The evaluation includes 9 image-language benchmarks, namely GQA [13], MMB [22], MME [5], POPE [17], SQA [24], TextVQA [30], OCRBench [23], SEEDBench [15], and OKVQA [26], as well as 4 video-language benchmarks, namely EgoSchema [25], NExTQA [36], VideoMME [10], and EgoPlan [7]. For base VLMs, we choose image-based models including LLaVA-1.5-7B [18], LLaVA-1.5-13B [18], LLaVA-Next-7B [19], and Qwen2-VL-7B [33], together with the video-based model LLaVA-Video-7B [43]. Please refer to the supplementary materials for more details.
@@ 4.2 Implementation Details
For LLaVA-1.5 [20] with a fixed number of visual tokens, we set $H=W=24$ ($N=576$), so that $D_{\max}=24\sqrt{2}$, which exactly matches the normalization in Eq. (5). For LLaVA-Next [19] and Qwen2-VL [33], which adopt a dynamic number of tokens depending on the image resolution, $D_{\max}$ is computed according to the actual height and width of the token feature map. For LLaVA-Video [43], which involves three spatial-temporal dimensions (height, width, and frames), the computation is correspondingly performed in the 3D coordinate space. Meanwhile, we set $\lambda=0.5$, $\tau^{(0)}=0.8$, $\Delta\tau=0.1$, $q=256$, $\beta=0.3$, and $B=16$.
All experiments are performed on a setup consisting of 8 GPUs, each with 32 GB of VRAM. We adhere to the official evaluation protocols established for both LLaVA-1.5 and LLaVA-Next. Additionally, for evaluating Qwen2-VL and LLaVA-Video, we utilize the lmms-evals package [41]. All results are obtained with a batch size of 1.
@@ 4.3 Comparative Experiment
We select methods based on importance-based baselines, such as FastV [6], SparseVLM [42], PDrop [37], and MustDrop [21], as well as methods based on redundancy-based baselines, such as DART [35] and DivPrune [2], as the comprehensive baselines for comparison.
@@ 4.4 Efficiency Analysis
We analyze the efficiency of VLM-Pruner on LLaVA-1.5-7B and Qwen2-VL-7B. Specifically, we report FLOPs following [6] and total inference time on the POPE and OCRBench benchmarks under a $\downarrow$88.9% pruning rate (i.e., retaining 11.1% of visual tokens), as shown in Table 5. VLM-Pruner outperforms the importance-based baselines SparseVLM and MustDrop in both speedup and FLOPs. In comparisons with redundancy-based methods, DivPrune exhibits slightly fewer FLOPs. This is because it prunes at layer 0, while VLM-Pruner and DART start at layer 2. However, this DivPrune setting comes at the cost of significantly poorer performance compared to our method (Sec. 4.3). VLM-Pruner’s inference speed matches DART. On Qwen2-VL-7B, which has more raw visual tokens, VLM-Pruner excels in detail-intensive tasks like OCRBench. It achieves the fastest speed (1.60×) and the best performance (Table 3). These results demonstrate VLM-Pruner’s superior capability for handling high-resolution inputs.
@@ 7 Details for Benchmarks
We select a broad range of commonly used tasks and benchmarks. Specifically, the evaluation includes 9 image-language benchmarks, namely GQA [13], MMB [22], MME [5], POPE [17], SQA [24], TextVQA [30], OCRBench [23], SEEDBench [15], and OKVQA [26], as well as 4 video-language benchmarks, namely EgoSchema [25], NExTQA [36], VideoMME [10], and EgoPlan [7].